Glenn Dental Office provides general dental care to residents of Philadelphia on a walk-in basis. The clinic has started receiving complaints from patients that the waiting time is too long and has asked you to investigate whether this problem can be solved. Upon arrival, customers first receive a series of paperwork from the receptionist and fill out relevant information such as personal health records and insurance provider. The form is then handed back to a receptionist who enters the information into the computer system for the dentist to see. A dental assistant then takes an X-ray from the patient. A dentist then performs the check-up and discusses any issues with the patient. Based on conversations with staff members at the clinic, you have obtained the following information on the process:

  • It takes about 3 minutes for a customer to fill out the paperwork.  

  • Entry of information on the paperwork into the system and verification with past records takes another 8 minutes for a receptionist. There are two receptionists.

  • It takes 15 minutes on average for the dental assistant to take an X-ray. There are six dental assistants on shift at any moment.       

  • There are eight dentists working at the clinic. Each check-up takes 30 minutes on average.

The following table summarizes the process data collected above.

Resource

Process

Staffing

(Numbers of Workers)

Activity Time

(Minutes per Patient)

Self-Service

Paperwork

-

3

Receptionists

Data Entry

2

8

Dental Assistant

X-ray

6

15

Dentist

Checkup

8

30

Assume that there exists unlimited demand, unless otherwise stated. Assume further that the dentist gets paid $120 per hour while receptionist and dental assistant get paid $30 per hour.             

f. Assume the process started empty, how long (in minutes) would it take 10 patients to go through the system? Round your answer to the nearest integer in minutes. For example, if your answer is 123.45 minutes, input 123.

Your answer is

.

g. What is the labor content in minutes? Round your answer to the nearest integer in minutes. For example, if your answer is 123.45 minutes, input 123.

Your answer is

.

h. What is the cost of direct labor? Round your answer to the nearest integer in minutes. For example, if your answer is 123.45 minutes, input 123.

Your answer is

.
Answer #1

Answer g- Total time for the production of N units=Sum of processing times + (N − 1)/1 / Flow rate

Flow rate is equal to the capacity of bottleneck resource Receptionists =2*60/8=15 patients / hr,

N=10

Total time for the production of 10 patients =(3+8+15+30)+(10-1)*60/15=92 min

Anwer h=Labor content =(2*8+6*15+8*30)=346 min

Answer i= cost of direct labor = (2*8*30/60)+(6*15*30/60)+(8*30*120/60) =$533